OpenStudy (anonymous):

Determine the equation of the line that passes through the points (2, -5) and (7, 0). Use the form y = mx + b. PLEASE HELP

OpenStudy (anonymous):

-5=m2+b 0=m7+b (-5-b)/2=-b/7 -35-7=-2b -42=-2b -21=-b

OpenStudy (anonymous):

-5=m2+21 -26=-m2 -13=-m

OpenStudy (anonymous):

y=13x+21

OpenStudy (anonymous):

http://www.purplemath.com/modules/strtlneq.htm Scroll down to "Find the equation of the line that passes through the points (–2, 4) and (1, 2)."